OpenStudy (anonymous):

Can anyone please help me answer this qestions please!!! The area of a rectangle is 24 square meters. Find the length and width of the rectangle if its length is 5 meters greater than its width. Use an equation and the formula for Area of a rectangle = (width)(length

OpenStudy (anonymous):

8 and 3

OpenStudy (anonymous):

width = 2.1908902300206644538278791312032 length = 10.954451150103322269139395656016

OpenStudy (anonymous):

area = 2.1908902300206644538278791312032 * 10.954451150103322269139395656016 = 24

OpenStudy (anonymous):

let width = x then length = 5x area = x * 5x 24 = 5x^2 .............. .........